# The heat flow on glued spaces with varying dimension

### v1

@@ 2.1 Basic theorems for $A_{2}-$weights on manifolds
In this section, we will show that, if we have a smooth compact manifold $M$ and a Muckenhoupt weight $\omega\in A_{2}(M)$, then the measure $\mu=\omega\cdot\text{vol}_{M}$ is doubling and fulfills the $2-$Poincare inequality and a reverse Hölder-inequality. These theorems are mostly due to [36, Chapter 15] where they are carried out in the case of a flat Euclidean space.
@@
Muckenhoupt weights satisfy a volume doubling condition.
@@
For $E\subset B$ where $B$ is a ball we have:
$\text{vol}(E)$ $=\int\limits_{E}\frac{\omega^{\frac{1}{2}}}{\omega^{\frac{1}{2}}}\mathop{}\!\mathrm{d}x$
@@
$\Rightarrow\mu(B)$ $\lesssim\left(\frac{\text{vol}(B)}{\text{vol}(E)}\right)^{2}\mu(E).$
A direct consequence is the volume doubling for $E=B_{r},B=B_{2r}$ as the volume measure of a compact manifold is doubling:
@@
The balls $B_{r}^{M}$ on $M$ are transformed by transition map $\phi$ to $A_{r}$. Since $\phi$ can be chosen Lipschitz, there exists $C>1$ s.t. $B_{\frac{r}{C}}\subset A_{r}\subset B_{Cr}$ and thus (since the metric is bounded)
$\text{vol}(A_{r})\simeq|A_{r}|\simeq|B_{r}|.$
@@
We conclude that $\omega$ (to be concrete $\omega\circ\phi^{-1}$) is also a Muckenhoupt weight on the flat chart, similar for $\omega\in A_{p}$, see Section 5. Hence, we can follow the usual proofs of [36, Chapter 15] to get a $p-$Poincare inequality without dilation. First, we can prove the doubling condition and also the reverse doubling condition which states for each ball $B$ and measurable subset $E\subseteq B$ there exists $10<q<$ 1 s.t.
@@
From this, we directly conclude the reverse Hölder inequality: There is $r>1$ s.t.
@@
This inequality can be used to obtain the open end property, i.e. if $\omega\in A_{p}$ with $p>1$ then $\omega\in A_{p-\varepsilon}$ for some $\varepsilon>0$. Concluding, this yields an equivalent characterization via the continuity of the Hardi-Littlewood maximal function on $L^{p}(\mu)-$functions, which finally proves a Poincare inequality on local balls for $\mu=\omega\cdot\text{vol}_{M}$ with $\omega\in A_{2}$:
$\forall r<R:\int\limits_{B_{r}}|f-(f)_{B_{r}}|^{2}\mathop{}\!\mathrm{d}\mu\lesssim r^{2}\int\limits_{B_{r}}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu.$
@@ 3 Proof of main theorem
3.1 Definitions
Given the setting of Section 2, we will prove our main theorem, Theorem 2.2. This theorem can be extended into the more general case, for this see Section 4.
The proof consists of three different steps to prove properties of the resulting space as well as of the operators and semi group on it. This will be done in the following. First, we define the relevant operators and function spaces. Afterwards, we will prove a continuity in a capacity sense. This, we will use to show a Rellich-Kondrachov embedding which yields ergodicity and irreducibility.
@@ 3.1 Definitions
We define the energy norm
$\left\lVert f\right\rVert_{H^{1}_{*}}:=\sqrt{\left\lVert f\right\rVert_{L^{2}(X,\mu)}^{2}+\operatorname*{\mathcal{E}}(f)}$
on the domain of $\mathcal{E}$
$D(\operatorname*{\mathcal{E}}):=\overline{C^{\infty}(\mathbb{R}^{d})}^{\left\lVert\cdot\right\rVert_{H^{1}_{*}}}=:H^{1}_{*}.$
Here, $d$ is the dimension of the surrounding space where $X$ is embedded in. The asterisk symbolizes the continuity which we will see in Section 3.2. Heuristically, this can be seen as a pointwise continuity on the intersection and Neumann conditions at the boundary as well as flow conditions at points where the manifolds meet.
Additionally, we define the Laplace-Beltrami operator $\operatorname*{\mathcal{A}}$ with its domain $D(\operatorname*{\mathcal{A}})$ via the Dirichlet form in the following way
$D(\operatorname*{\mathcal{A}}):=\{f\in D(\operatorname*{\mathcal{E}})|\exists\operatorname*{\mathcal{A}}f\in L^{2}(X,\mu):(\operatorname*{\mathcal{A}}f,g)=-\operatorname*{\mathcal{E}}(f,g)\;\forall g\in D(\operatorname*{\mathcal{E}})\}.$
By definition, this is a selfadjoint symmetric non-positive operator which induces an analytic semigroup $S_{h}$ that we call the heat semigroup. The properties of $\operatorname*{\mathcal{A}}$ are computed via the resolvent
@@
In Section 3.3, we will show and use that $R_{\operatorname*{\mathcal{A}}}$ is compact to prove a spectral gap on $\operatorname*{\mathcal{A}}$.
@@ 3.2 Capacity bounds
Under the assumptions of Theorem 2.2, the capacity of the intersection is positive in both manifolds:
$\text{Cap}_{2,\mu_{1}}(L,X)>0,\quad\text{Cap}_{2,\mu_{2}}(L,X)>0.$
Proof.
We recall that intersection of the manifolds $L$ is a smooth manifold of dimension $k$.
In the following, we treat the intersection as a linear subspace $L\subseteq\mathbb{R}^{k}$. This is rectified by considering charts of the manifolds s.t. they agree on the intersection $L$ and possess smooth (bi-Lipschitz) transition maps which map $L$ into the linear subspace. By compactness of $L$ and $M_{i}$, finitely many suffice and the capacity of $L$ in $M_{i}$ is finite iff there is one chart on which the relative capacity is finite.
Moreover, by continuity of the metric, the charts are chosen s.t. $\sqrt{\det g}$ is uniformly bounded from above and below and s.t. they agree on $L$ for both manifolds. Thus, we will work with the Lebesgue measure $\mathcal{L}^{n}$ instead of the volume measure $\text{vol}_{M}$ in this section as the calculations are equivalent up to a factor.
First, we will need an equivalence of the partial measures close to the intersection.
Let $R^{*}(x)$ be small enough s.t. for every $x^{*}\in L,R<R^{*}$ there is a point $x\in\partial B_{\frac{R}{2}}(x^{*})\cap M_{1}$ with $B_{\frac{R}{2}}(x)\subset M_{1}$. Then, we can estimate the partial measure $\mu_{i}$ on balls on the intersection. W.l.o.g. only written down in the direction $M_{2}\leq M_{1}$:
$\mu_{2}(B_{R}(x^{*}))$ $\leq\mu(B_{R}(x^{*}))\leq\mu(B_{\frac{3}{2}R}(x))\leq N(R)\mu(B_{\frac{R}{2}}(x)))$
$=N(R)\mu_{1}(B_{\frac{R}{2}}(x))\leq N(R)\mu_{1}(B_{R}(x^{*})).$
@@
$R^{*}(x)$ can be bounded uniformly from below since $M_{i}$ and thus also $L$ are compact manifolds.
The point $x$ needs to exist by dimensional arguments as $k+1=n_{2}<n_{1}$ and thus the dimension $d^{\prime}$ of the manifold in which $M_{1}$ and $M_{2}$ are transversal is larger than $n_{1}$ and $n_{2}$ as $d^{\prime}=n_{1}+n_{2}-k$. Therefore, there exists a direction solely in $T_{x^{*}}M_{1}$ for each $x^{*}\in L$ which yields existence of $x$ by smoothness of the manifolds.
To compute the capacity, it proves useful to consider a thickened version of the intersection $L_{R}:=\bigcup\limits_{x\in L}B_{R}(x)$ as we will estimate the relative capacity $\text{Cap}_{2,\mu}(L_{r},L_{R})$ on $L_{R}$ via an approximation $L_{r}\to L$ as $r\searrow 0$. In the following, we will estimate the capacity from below on both manifolds in order to ensure that the manifolds are sufficiently connected. The techniques are based on [36, Chapter 2] where the case of balls is proven.
For this, we need a representation formula.
In this, we will work on $\mathbb{R}^{n}$ by the estimate of the volume measure and the smoothness of the transition map this is justified as the calculations carry over to the charts and manifolds. Take $u\in C^{1}$ s.t. $u\equiv 0$ outside a compact set. Then, for a direction $w\in\mathbb{R}^{n}$ s.t. $|w|=1$:
@@
The procedure can be seen geometrically as cylinder like coordinates, see Figure 2. By integrating over the unit sphere in $n-k$ dimensions (extended with 00 components in $\mathbb{R}^{n}$) and with $x=0$ and $u(0)=1$ we get:
$c_{n-k}$ $=-\int\limits_{0}^{\infty}\int\limits_{\mathbb{S}^{n-k-1}}\partial_{r}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n-k-1}(w)\mathop{}\!\mathrm{d}r$
$=-\int\limits_{\mathbb{R}^{n-k}}r^{-(n-k-1)}\partial_{r}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(r\cdot w)$
$=-\int\limits_{\mathbb{R}^{n-k}}\frac{r}{|rw|^{n-k}}w\cdot\nabla^{n-k}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(r\cdot w)$
$=-\int\limits_{\mathbb{R}^{n-k}}\frac{y\cdot\nabla^{n-k}u(y)}{|y|^{n-k}}\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(y).$
@@
11 11 $\simeq\left(\int\limits_{L}\mathop{}\!\mathrm{d}\mathcal{H}^{k}(x)\right)^{2}$
$\simeq\left(\int\limits_{L}\int\limits_{\mathbb{R}^{n-k}(x)}\frac{(x-y)\cdot\nabla^{n-k}u(y)}{|x-y|^{n-k}}\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(y)\mathop{}\!\mathrm{d}\mathcal{H}^{k}(x)\right)^{2}$
$\lesssim\int\limits_{L_{R}}|\nabla^{n-k}u|^{2}\mathop{}\!\mathrm{d}\mu_{i}\cdot\int\limits_{L_{R}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}(y)}{\omega_{i}(y)}.$
@@
$\text{Cap}_{2,\mu_{i}}(L,L_{R})$ $=\text{Cap}_{2,\mu_{i}}(\bigcap\limits_{r>0}L_{r},L_{R})=\lim\limits_{r\searrow 0}\text{Cap}_{2,\mu_{i}}(L_{r},L_{R})$
$\geq\lim\limits_{r\searrow 0}\inf\left\{\int\limits_{L_{R}}|\nabla^{n-k}u|^{2}\mathop{}\!\mathrm{d}\mu_{i}:u\geq\mathds{1}_{L_{r}}\right\}$
$\gtrsim\left(\int\limits_{L_{R}}\frac{1}{\operatorname*{dist}(y,L)^{2(n-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}(y)}{\omega_{i}(y)}\right)^{-1}.$
[…]
Figure 2: We prove the lower capacity bound via a foliation procedure.
Next, we note that in our setting the density $\omega_{2}$ is in such a way that
$\int_{L_{r}}\omega_{2}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\int_{L_{r}}\frac{1}{\omega_{2}}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\simeq r^{2(n_{2}-k)}$
as it is bounded from above and below. This can be compared to a Muckenhoupt condition on the level of $L$.
The lower bound can now be simplified into a more suitable version and connected to the other manifold. By the layer cake formula and the transformation $\rho=t^{-\frac{1}{2(n-k-1)}}$:
$\int\limits_{L_{R}}$ $\frac{1}{\operatorname*{dist}(y,L)^{2(n-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n}}{\omega}$
$=\int\limits_{0}^{\infty}\frac{\mathcal{L}^{n}}{\omega}(x\in L_{R}:\operatorname*{dist}(x,L)^{-2(n-k-1)}>t)\mathop{}\!\mathrm{d}t$
$=R^{-2(n-k-1)}\frac{\mathcal{L}^{n}}{\omega}(L_{R})+\int\limits_{R^{-2(n-k-1)}}^{\infty}\frac{\mathcal{L}^{n}}{\omega}(L_{t^{-\frac{1}{2(n-k-1)}}})\mathop{}\!\mathrm{d}t$
$=R^{-2(n-k-1)}\frac{\mathcal{L}^{n}}{\omega}(L_{R})+2(n-k-1)\int\limits_{0}^{R}\rho^{-(2n-2k-1)}\frac{\mathcal{L}^{n}}{\omega}(L_{\rho})\mathop{}\!\mathrm{d}\rho.$
To combine the estimates on both manifolds, we now want to compare $\frac{\mathcal{L}^{n}}{\omega}(L_{R})$ (and $\omega\mathcal{L}^{n}(L_{R})$) on both manifolds to get an estimate of $\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(L_{R})$ by $\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})$. To this end, consider the Vitali covering theorem with the covering of $L_{R}$ by $B_{R}(x)$ for $x\in L$. There exist finitely many $x_{j}\in L$ s.t. $L_{R}\subset\bigcup 5B_{j}$ ($5B_{j}:=B_{5R}(x_{j})$) and $B_{R}(x_{j})$ are disjoint. Since $\omega$ is a 2-Muckenhoupt weight, we know that $\omega\mathcal{L}^{n}(B_{R})$ and $\frac{\mathcal{L}^{n}}{\omega}(B_{R})$ can be compared: $\omega\mathcal{L}^{n}(B_{R})\cdot\frac{\mathcal{L}^{n}}{\omega}(B_{R})\simeq R^{2n}$.
@@
Together, this yields:
@@
$\gtrsim\left(c_{R}+c\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
@@
$\text{Cap}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+c\int\limits_{0}^{R}N(\rho)\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
$\text{Cap}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
@@
as we fixed $R>0$ and $\mu_{2}(L_{\rho})\simeq\rho$. The dimensional condition on $k$ is necessary for $\mu_{2}(L)=0$ and $\text{Cap}_{2,\mu_{2}}(L,L_{R})>0$.
Thus, the intersection has positive capacity in $M_{1}$ in this setting if the same holds for $M_{2}$ and $\mu$ satisfies a $N$-doubling condition with integrable $N$.
Since the global and restricted capacity are equivalent for fixed $R>0$, this implies that q.e. point of the intersection is a Lebesgue point, c.f. [40, Theorem 3.5] and [41, Corollary 3.7]. Therefore, one has a trace in the sense of a quasi-continuous refinement on $L$. ∎

### v2

@@ 2.1 Basic Theorems for $A_{2}$-Weights on Manifolds
In this section, we show that, if we have a smooth compact manifold $M$ and a Muckenhoupt weight $\omega\in A_{2}(M)$, then the measure $\mu=\omega\cdot{\text{vol}}_{M}$ is doubling and fulfills the $2$-Poincaré inequality and a reverse Hölder inequality. These theorems are mostly due to [40, Chapter 15] where they are carried out in the case of a flat Euclidean space. But one can also work in charts to obtain them on smooth manifolds.
@@
A Muckenhoupt weight satisfies a volume doubling condition:
$\mu(B_{2r})\lesssim\mu(B_{r})\quad\forall r>0.$
@@ A Muckenhoupt weight satisfies a volume doubling condition:
For $E\subseteq B$ where $B$ is a ball, we have:
$\text{vol}_{M}(E)$ $=\int\limits_{E}\frac{\omega^{\frac{1}{2}}}{\omega^{\frac{1}{2}}}\mathop{}\!\mathrm{d}x$
@@
$\Rightarrow\mu(B)$ $\lesssim\left(\frac{\text{vol}_{M}(B)}{\text{vol}_{M}(E)}\right)^{2}\mu(E).$
A direct consequence is the volume doubling for $E=B_{r},B=B_{2r}$ since the volume measure of a compact manifold is doubling:
@@
For this, take balls $B_{r}^{M}$ in $M$. They are transformed by transition map $\phi$ to $A_{r}$. Since $\phi$ can be chosen to be Lipschitz, there exists $c\geq 1$ such that $B_{\frac{r}{c}}\subseteq A_{r}\subseteq B_{cr}$ and thus, since the metric is bounded,
$\text{vol}_{M}(B_{r}^{M})\simeq|A_{r}|\simeq|B_{r}|.$
@@
We conclude that $\omega$ (to be precise, $\omega\circ\phi^{-1}$) is also a Muckenhoupt weight on the flat chart, similar for $\omega\in A_{p}$, see Section 5. Hence, we can follow the usual proofs of [40, Chapter 15] to obtain a $p$-Poincaré inequality without dilation, cf. Definition 5.2. First, we can prove the doubling condition and also the reverse doubling condition which states that for each ball $B$ and measurable subset $E\subseteq B$ there exists $0<q<1$ such that
@@
From this, we directly conclude the reverse Hölder inequality: There is $r>1$ such that
@@
This inequality can be used to obtain the open end property, that is, if $\omega\in A_{p}$ with $p>1$ then $\omega\in A_{p-\varepsilon}$ for some $\varepsilon>0$. Concluding, this yields an equivalent characterization via the continuity of the Hardy-Littlewood maximal function on $L^{p}(\mu)$-functions, which finally proves a Poincaré inequality on local balls for $\mu=\omega\cdot{\text{vol}}_{M}$ with $\omega\in A_{2}(M)$:
$\int\limits_{B_{r}}|f-f_{B_{r}}|^{2}\mathop{}\!\mathrm{d}\mu\lesssim r^{2}\int\limits_{B_{r}}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu.$
It is noteworthy that all consideration could also be done from the perspective of Brownian motion instead of the heat flow as they are equivalent.
In this paper, we focus on the perspective of the heat flow. The Brownian motion is a Markov process with the Laplace-Beltrami operator as infinitesimal generator. More general, we define the Brownian motion on a locally compact metric measure space with strongly local Dirichlet form as the Markov process with the semigroup $S_{h}$. This semigroup is constructed from the Laplace operator which comes from the Dirichlet form as explained in this and the next section; cf. [24].
The heat flow and the Brownian motion are closely related to each other. In $\mathbb{R}^{n}$ or smooth manifolds, one could think of the heat kernel as the transition density of the Brownian motion. Similarly, we can solve the heat flow using the Brownian motion as the evaluation points, i.e., $u(x,t)=\mathbb{E}_{x}[u(B_{t},0)]$. An example of a sample path of the Brownian motion on the sphere can be seen in Figure 2.
Figure 2: A sample path of the Brownian motion on a sphere.
@@ 3 Proof of Main Theorem
3.1 Definition of the Heat Flow
Given the setting of Section 2, we prove our main theorem, Theorem 2.5. This theorem can be extended to the more general case; for this, see Section 4. This includes the extension to multiple manifolds glued together with general Muckenhoupt weights on them.
The proof consists of three different steps to prove properties of the resulting space as well as of the operators and semigroup on it. This is done in the following. First, we define the relevant operators and function spaces. Afterwards, we prove capacity bounds. This we use to show a Rellich-Kondrachov embedding which yields ergodicity and irreducibility.
@@ 3.1 Definition of the Heat Flow
We recall from Section 2 the energy norm
$\left\lVert f\right\rVert_{H^{1}_{*}}\coloneqq\sqrt{\left\lVert f\right\rVert_{L^{2}(X,\mu)}^{2}+\operatorname*{\mathcal{E}}(f)}$
and the domain of $\mathcal{E}$
$D(\operatorname*{\mathcal{E}})\coloneqq\overline{C^{\infty}(\mathbb{R}^{\widetilde{n}})}^{\left\lVert\cdot\right\rVert_{H^{1}_{*}}}=:H^{1}_{*}.$
Here, $\widetilde{n}$ is the dimension of the surrounding space $X$ is embedded in and we associate functions if their difference has norm zero. The star symbolizes the continuity which we see in Section 3.2. Heuristically, this can be seen as a pointwise continuity on the intersection and Neumann conditions at the boundary as well as flow conditions at points where the manifolds meet. As discussed in Section 2, one can also choose an intrinsic definition via $\operatorname*{Lip}_{\operatorname*{\mathcal{E}}}$.
Additionally, we define the Laplace-Beltrami operator $\operatorname*{\mathcal{A}}$ with its domain $D(\operatorname*{\mathcal{A}})$ using the Dirichlet form in the following way
$D(\operatorname*{\mathcal{A}})\coloneqq\{f\in D(\operatorname*{\mathcal{E}})|\exists\operatorname*{\mathcal{A}}f\in L^{2}(X,\mu):(\operatorname*{\mathcal{A}}f,g)=-\operatorname*{\mathcal{E}}(f,g)\;\forall g\in D(\operatorname*{\mathcal{E}})\}.$
[…]
@@
In Section 3.3, we show and use that $R_{\operatorname*{\mathcal{A}}}$ is compact to prove a spectral gap on $\operatorname*{\mathcal{A}}$.
@@ 3.2 Capacity Bounds
Given the assumptions of Theorem 2.5, the capacity of the intersection is positive in both manifolds, that is, there exists $R>0$ such that
${\text{Cap}}_{2,\mu_{1}}(L,L_{R})>0,\quad{\text{Cap}}_{2,\mu_{2}}(L,L_{R})>0$
where $L_{R}\coloneqq L+B_{R}=\bigcup\limits_{x\in L}B_{R}(x)$.
The proof is given in this subsection. We recall that the intersection of the manifolds $L$ is a smooth manifold of dimension $k$.
In the following, we treat the intersection as a linear subspace $L\subseteq\mathbb{R}^{k}$. This is justified by considering charts of the manifolds such that they agree on the intersection $L$ and possess smooth (bi-Lipschitz) transition maps that map $L$ into the linear subspace. By compactness of $L$ and $M_{i}$, finitely many suffice and the capacity of $L$ in $M_{i}$ is finite iff there is a chart in which the relative capacity is finite.
Moreover, by the continuity of the metric, the charts are chosen such that $\sqrt{\det g}$ is uniformly bounded from above and below and so that they agree on $L$ for both manifolds. Thus, in this section, we work with the Lebesgue measure $\mathcal{L}^{n}=\mathop{}\!\mathrm{d}x$ instead of the volume measure ${\text{vol}}_{M}$ because the calculations are equivalent up to a factor.
First, we need an equivalence of the partial measures close to the intersection.
Let $R_{*}(x)$ be small enough so that for every $x_{*}\in L,R<R_{*}$ there is a point $x\in\partial B_{\frac{R}{2}}(x_{*})\cap M_{1}$ with $B_{\frac{R}{2}}(x)\subseteq M_{1}$. Then, we can estimate the partial measure $\mu_{i}$ on the balls at the intersection. Without loss of generality, only written down in the direction $M_{2}\leq M_{1}"$:
$\mu_{2}(B_{R}(x_{*}))$ $\leq\mu(B_{R}(x_{*}))\leq\mu(B_{\frac{3}{2}R}(x))\leq N(R)\mu(B_{\frac{R}{2}}(x)))$
$=N(R)\mu_{1}(B_{\frac{R}{2}}(x))\leq N(R)\mu_{1}(B_{R}(x_{*})).$
@@
$R_{*}(x)$ can be uniformly bounded from below, since $M_{i}$ and thus also $L$ are compact manifolds.
The point $x$ must exist by dimensional arguments as $k<k+1=n_{2}<n_{1}$. Therefore, there exists a direction solely in $T_{x_{*}}M_{i}$ for each $x_{*}\in L$ that yields the existence of $x$ by the smoothness of the manifolds.
To compute the capacity, it proves useful to consider a thickened version of the intersection $L_{R}\coloneqq\bigcup\limits_{x\in L}B_{R}(x)$ as we estimate the relative capacity ${\text{Cap}}_{2,\mu}(L_{r},L_{R})$ on $L_{R}$ and use an approximation $L_{r}\to L$ as $r\searrow 0$. In the following, we estimate the capacity below for both manifolds to ensure that the manifolds are sufficiently connected. The techniques are based on [40, Chapter 2] where the case of balls is proven.
For this calculation, we need a representation formula. We work on $\mathbb{R}^{n_{i}}$. By the estimate of the volume measure and the smoothness of the transition map, this is justified, as the calculations carry over to the charts and manifolds. Fix a function $u\in C^{1}_{c}$. Then, for a direction $w\in\mathbb{R}^{n_{i}}$ such that $|w|=1$:
@@
The procedure can be seen geometrically as cylinder like coordinates, see Figure 3. By integrating over the unit sphere in dimension $n_{i}-k$ (extended with 0 components in $\mathbb{R}^{n_{i}}$) and with $x=0$ and $u(0)=1$, we get the following:
$c_{n_{i}-k}$ $=-\int\limits_{0}^{\infty}\int\limits_{\mathbb{S}^{n_{i}-k-1}}\partial_{r}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n_{i}-k-1}(w)\mathop{}\!\mathrm{d}r$
$=-\int\limits_{\mathbb{R}^{n_{i}-k}}r^{-(n_{i}-k-1)}\partial_{r}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n_{i}-k}(r\cdot w)$
$=-\int\limits_{\mathbb{R}^{n_{i}-k}}\frac{r}{|rw|^{n_{i}-k}}w\cdot\nabla^{n_{i}-k}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n_{i}-k}(r\cdot w)$
$=-\int\limits_{\mathbb{R}^{n_{i}-k}}\frac{y\cdot\nabla^{n_{i}-k}u(y)}{|y|^{n_{i}-k}}\mathop{}\!\mathrm{d}\mathcal{H}^{n_{i}-k}(y).$
@@
$1$ $\simeq\left(\int\limits_{L}\mathop{}\!\mathrm{d}\mathcal{H}^{k}(x)\right)^{2}$
$\simeq\left(\int\limits_{L}\int\limits_{\mathbb{R}^{n_{i}-k}(x)}\frac{(x-y)\cdot\nabla^{n_{i}-k}u(y)}{|x-y|^{n_{i}-k}}\mathop{}\!\mathrm{d}\mathcal{H}^{n_{i}-k}(y)\mathop{}\!\mathrm{d}\mathcal{H}^{k}(x)\right)^{2}$
$\lesssim\int\limits_{L_{R}}|\nabla^{n_{i}-k}u|^{2}\mathop{}\!\mathrm{d}\mu_{i}\cdot\int\limits_{L_{R}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}(y)}{\omega_{i}(y)}.$
@@
${\text{Cap}}_{2,\mu_{i}}(L,L_{R})$ $={\text{Cap}}_{2,\mu_{i}}\left(\bigcap\limits_{r>0}L_{r},L_{R}\right)=\lim\limits_{r\searrow 0}{\text{Cap}}_{2,\mu_{i}}(L_{r},L_{R})$
$\geq\lim\limits_{r\searrow 0}\inf\left\{\int\limits_{L_{R}}|\nabla^{n_{i}-k}u|^{2}\mathop{}\!\mathrm{d}\mu_{i}:u\geq\chi_{L_{r}}\right\}$
$\gtrsim\left(\int\limits_{L_{R}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}(y)}{\omega_{i}(y)}\right)^{-1}.$
$\mathbb{R}^{n_{i}-k}$$\mathbb{S}^{n_{i}-k}$$L\subseteq\mathbb{R}^{k}$
Figure 3: We prove the lower capacity bound via a foliation procedure as shown in this figure.
Next, we note that in our setting the density $\omega_{2}$ is chosen in such a way that a $L$-Muckenhoupt type condition on the intersection is satisfied.
Definition 3.2 ($L$-Muckenhoupt condition).
We say that $\mu=\omega\cdot{\text{vol}}_{M}$ satisfies the $L$-Muckenhoupt type condition for a submanifold $L$ in $M$ with $\text{dim}(L)=k,\text{dim}(M)=n$, if:
$\int_{L_{R}}\omega\mathop{}\!\mathrm{d}\text{vol}_{M}\int_{L_{R}}\frac{1}{\omega}\mathop{}\!\mathrm{d}\text{vol}_{M}\simeq R^{2(n-k)}.$
This is due to the fact that $\omega_{2}$ is bounded from above and below. This can be compared to a Muckenhoupt type condition on the level of $L$.
The lower bound can now be simplified into a more suitable version and connected to the other manifold. By the layer cake formula and the transformation $\rho=t^{-\frac{1}{2(n_{i}-k-1)}}$:
$\int\limits_{L_{R}}$ $\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}}{\omega}$
$=\int\limits_{0}^{\infty}\frac{\mathcal{L}^{n_{i}}}{\omega}(x\in L_{R}:\operatorname*{dist}(x,L)^{-2(n_{i}-k-1)}>t)\mathop{}\!\mathrm{d}t$
$=R^{-2(n_{i}-k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega}(L_{R})+\int\limits_{R^{-2(n_{i}-k-1)}}^{\infty}\frac{\mathcal{L}^{n_{i}}}{\omega}\left(L_{t^{-\frac{1}{2(n_{i}-k-1)}}}\right)\mathop{}\!\mathrm{d}t$
$=R^{-2(n_{i}-k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega}(L_{R})+2(n_{i}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{i}-2k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega}(L_{\rho})\mathop{}\!\mathrm{d}\rho.$
To combine the estimates on both manifolds, we now compare $\frac{\mathcal{L}^{n_{i}}}{\omega}(L_{R})$ on both manifolds to get an estimate of $\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(L_{R})$ by $\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})$. To this end, consider the Vitali covering theorem with the covering of $L_{R}$ by $B_{R}(x)$ for $x\in L$. There exist finitely many $x_{j}\in L$ such that $B_{R}(x_{j})$ are disjoint and
$L_{R}\subseteq\bigcup 5B_{j}$
where $5B_{j}\coloneqq B_{5R}(x_{j})$. Since $\omega_{i}$ is a 2-Muckenhoupt weight, we know that $\omega_{i}\mathcal{L}^{n_{i}}(B_{R})$ and $\frac{\mathcal{L}^{n_{i}}}{\omega_{i}}(B_{R})$ can be compared:
$\omega_{i}\mathcal{L}^{n_{i}}(B_{R})\cdot\frac{\mathcal{L}^{n_{i}}}{\omega_{i}}(B_{R})\simeq R^{2n_{i}}.$
@@
Together, this yields the following:
@@
$\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+c\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
@@
${\text{Cap}}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+c\int\limits_{0}^{R}N(\rho)\rho\mu_{2}(L_{\rho})^{-1}\mathop{}\!\mathrm{d}\rho\right)^{-1},$
${\text{Cap}}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}(L_{\rho})^{-1}\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
@@
as we fixed $R>0$ and $\mu_{2}(L_{\rho})\simeq\rho$. The dimensional condition on $k$ for a bounded weight is necessary for $\mu_{2}(L)=0$ and ${\text{Cap}}_{2,\mu_{2}}(L,L_{R})>0$.
Thus, the intersection has a finite positive capacity in $M_{1}$ in this setting if the same holds for $M_{2}$ and $\mu$ satisfies an $N$ double condition with integrable $N$.
Since the global and restricted capacity are equivalent for fixed $R>0$, this completes the proof.
Remark 3.
Moreover, this implies that q.e. point of the intersection is a Lebesgue point, cf. [44, Theorem 3.5] and [45, Corollary 3.7]. Therefore, we have a trace in the sense of a quasi-continuous refinement on $L$.